# EgoAgent: A Joint Predictive Agent Model in Egocentric Worlds

### v2

@@ 3.3 Task-Specific Inference
After training, EgoAgent inherently acquires the abilities to solve the three tasks designed to mirror human cognition.
World state prediction. Given a sequence of egocentric video clips and human motion, EgoAgent is capable of predicting the future world state $S_{t+1}$ by $S_{t+1}=\Theta(I_{1},A_{1},...,I_{t},A_{t}).$ Once predicted, the future world image can then be retrieved by measuring similarity within the learned feature space, enabling applications that require foresight into upcoming visual states.
Human action prediction. From an egocentric perspective and informed by previous actions, EgoAgent can generate plausible future action $A_{t}$ represented as 3D human motion by $A_{t}=\Theta(I_{1},A_{1},...,A_{t-1},I_{t}).$ This predictive capability of human motion is essential for applications in humanoid robotics, virtual environments, and interactive gaming, where anticipating human-like movements is crucial.
Visual representation. EgoAgent learns world states as feature representations, allowing it to extract meaningful features from input images by $S=\Theta(I),$ where $\Theta$ denotes the EgoAgent model. These features can be directly applied to representation tasks such as image classification.
Table 1: Performance comparison across three tasks. (1) For world state prediction, we report the feature retrieval accuracy of the predicted next-frame features by the averaged Top1 accuracy and mAP on total timesteps $T_{s}=4$. (2) For human action prediction, we report the 3D motion prediction results on Ego-Exo4D by MPJPE (cm) and MPJVE (cm/s) with time gaps of 1/30 and 1/10 seconds. (3) For visual representation, we report the $k$-NN evaluation results on ImageNet-100 and ImageNet-1K by Top1 and Top5 accuracy (%). We evaluate with the officially released checkpoints of representation models trained on egocentric video datasets and train all motion prediction models from scratch. denotes the model trained by ourselves for a fair comparison.
@@ 4.1 Experimental Setups
Datasets. We train EgoAgent on two egocentric datasets, WalkingTours (WT) [44] and Ego-Exo4D [20]. WT is a video-only dataset with around 1.5M high-resolution frames captured across cities worldwide. Ego-Exo4D has 221.26 hours of egocentric videos along with 376K manually labeled and 9.2M automatically generated 3D body poses. To eliminate projection differences between Ego-Exo4D and WT, we undistort the raw fisheye sensor data into a pinhole camera model. We also apply a 20-frame sliding window filter to extract continuous pose sequences, resulting in 1,410,119 clips of synchronized egocentric videos and 3D body poses. Based on the EgoPose split in Ego-Exo4D, we reserve the validation set of Ego-Exo4D-v2333The whole Ego-Exo4D dataset is divided into Ego-Exo4D-v1 and Ego-Exo4D-v2 on the official website (https://ego-exo4d-data.org). for evaluation and use the remaining clips for training, yielding 1,378,672 training clips and 31,447 evaluation clips.
Implementation Details. To alleviate the training burden, we sample one image frame every five frames while retaining all 3D body poses. Specifically, for each clip with 20 frames, we divide it into $T_{s}=4$ time steps and sample the first image frame from each step. This results in every time step consisting of one image frame paired with five frames of 3D body poses. We follow [44] to avoid cropping noisy positive pairs and adopt two global crops, six local crops, and the same augmentations consistent with DINO [10].
Following the configurations in [21], we train EgoAgent with two model sizes, 300M and 1B. We train EgoAgent on the processed dataset for 72,000 iterations with a linear warm-up of 1800 iterations. We set the base learning rate to $6\times 10^{-4}$ with a cosine decay scheduler. The loss weights are set as $\lambda_{rep}=2,\lambda_{pred}=1,\lambda_{act}=3$, respectively. All models are trained from scratch using FP16 to speed up. The training takes 25/60 hours in total with a batch size of 1920 on 32/48 NVIDIA A100 GPUs for EgoAgent-300M/EgoAgent-1B, respectively.
Figure 3: Retrieval results for egocentric world state prediction. Correctly retrieved images are marked with green boundaries.
@@ 4.2 World State Prediction
To evaluate the effectiveness of EgoAgent in predicting future states of the world, we perform an egocentric video prediction task with feature retrieval. Specifically, given the egocentric video frames and 3D human pose sequences at the current time step, we build a query set that stores the predicted features of the next frame and a gallery set that contains the features directly extracted from all video frames. At time step $t$, if the predicted world state $S_{t+1}^{\prime}$ in the query set can correctly retrieve the corresponding $S_{t+1}$ in the gallery set, we treat it as a successful prediction. For representation models, since there are no actions to condition the input image features, the predicted next state $S_{t+1}^{\prime}$ is exactly the state $S_{t}$ at time step $t$. To avoid retrieving this exact same state, when querying with input image $I_{t}$, we ignore the scene state $S_{t}$ in the gallery set. Following common practice in retrieval tasks, we use Top1 accuracy and mean average precision (mAP) as metrics.
As shown in Tab. 1, EgoAgent outperforms representation models by a large margin. Specifically, EgoAgent-300M improves the performance of DoRA by +12.86% in Top1 accuracy and +13.05% in mAP, indicating that it predicts the future states based not only on the semantic similarity of images. In contrast, representation models rely solely on semantic similarity, struggling with visually similar frames in egocentric videos. Furthermore, scaling up the model enhances performance, as EgoAgent-1B achieves further performance gains upon EgoAgent-300M with +3.42% Top1 accuracy and +3.90% mAP. Note that DoRA, when trained on our combined dataset, exhibits performance degradation in both world state prediction and visual representation tasks. We attribute this to its tracker’s inability to track abrupt object motions in Ego-Exo4D. For further analysis, please refer to Sec. 4.4. Fig. 3 visualizes the retrieval image by predicted features, where DoRA fails to retrieve the correct future states when humans have large movements, while EgoAgent successfully predicts the future state features that retrieve the real next-frame image.
@@
To further evaluate whether EgoAgent accurately models the causal relationship between actions and future world states, we conduct a qualitative analysis by conditioning next-state predictions on various poses with the same initial observing image, including dribbling, raising, shooting, and passing the ball. Fig. 4 shows the retrieved images, where all predicted features successfully retrieve images that accurately reflect the intended motion dynamics. This demonstrates EgoAgent’s ability to understand both the semantics of actions and their consequences in the egocentric world.
Figure 5: Visualizations of the 3D motion prediction tasks. After observing the egocentric frames, EgoAgent can generate accurate human skeletons even though most of the 3D body points are not visible in the input frame.
@@
Quantitatively, as shown in Tab. 1, EgoAgent achieves state-of-the-art performance with the lowest MPJPE and competitive MPJVE. Specifically, EgoAgent-300M improves siMLPe [22] by -0.41 and -0.31 MPJPE on the evaluation set of 30fps and 10 fps, respectively. When scaling up to 1B, the prediction errors are further reduced. These results demonstrate that EgoAgent can predict actions accurately when given egocentric observations and past actions.
We also qualitatively evaluate the motion prediction results. As illustrated in Fig. 5, EgoAgent produces relatively small errors compared to the ground truth. Compared to Diffusion Policy [13], EgoAgent achieves higher accuracy on non-visible joints. Specifically, in the early frames ($t_{1}$ and $t_{2}$), the missing right hand in the image leads to large prediction errors in Diffusion Policy, whereas EgoAgent accurately estimates hand positions by leveraging prior motion cues. Compared to HumanMAC [11], EgoAgent shows smaller accelerated errors. The predicted skeletons in HumanMAC gradually lean toward the floor ($t_{[3:5]}$) after several frames, whereas EgoAgent integrates visible body part information from egocentric observations to correct accelerated errors, leading to more stable motion predictions.
@@ 4.4 Visual Representation
We examine EgoAgent’s visual perception ability by two aspects: (1) image classification accuracy by $k$-NN on ImageNet-100 and ImageNet-1K [15], which is the golden standard for evaluating visual representations, and (2) egocentric manipulation success rate in an embodied simulator called TriFinger [48], which measures the transferability of learned visual representations to visual-motor control. Concretely, we freeze the model and extract features of images across two tasks, for image classification, we utilize a $k$-nearest neighbor classifier with $k=20$.
@@ 4.5 Ablation Study
Joint learning of representation, prediction, and action. EgoAgent demonstrates promising performance across the three tasks. To understand how these tasks interact during training, we conduct leave-one-out experiments and one-task-only experiments, as shown in Tab. 3 (a-g). We observe several key findings. First, removing any one task degrades the performance of the other two, highlighting the complementary roles of these tasks in learning. Specifically, removing the representation loss causes the largest drop in state prediction (-11.87% in Top1 accuracy and -13.55% in mAP), while removing the prediction loss significantly weakens representation learning, reducing Top-1 accuracy by -2.52% on ImageNet-100 and -1.31% on ImageNet-1K. This suggests that state prediction and visual representation are closely interconnected. For motion prediction, excluding the prediction loss leads to the most significant decline, increasing MPJVE by 1.01. Second, training on individual tasks yields worse performances on each task than learning all three tasks together, except for MPJPE in motion prediction. We suspect that when trained only on motion prediction, the model focuses excessively on minimizing joint position errors. However, without context from the other tasks, it struggles to maintain smooth, consistent motion across frames, resulting in a higher MPJVE of 2.52.
@@
Analysis of task dependencies. We also identify interesting dependencies between the three tasks. The key interactions are that actions rely on the understanding (or representation) of the current world state, while next-state predictions depend on both the current state understanding and actions taken. Specifically, (1) Representation serves as the foundation for both prediction and action as it captures semantic information to model the current world state. (2) Prediction and action jointly provide effective supervision for representation by modeling world dynamics and predicting actions, improving representation by 1.16% in Top1 accuracy. However, training on prediction or action alone cannot enhance representation, as actions without accurate next-state predictions or next-state predictions conditioned on incorrect actions may provide noisy supervision. (3) Representation and action guide the model to integrate learned representations with action dynamics, leading to more accurate predictions of future states. Meanwhile, representation and prediction work together to help the model understand how actions influence state transitions, refining its ability to generate actions.
@@
Training in semantic feature space. To evaluate the effectiveness of training agent models with a high-level semantic feature space, we use a pretrained VQGAN [17] as the image tokenizer. The results are shown in Tab. 3 (a,h,i). Training EgoAgent in VQGAN’s pretrained pixel-level latent space results in significant performance drops in world state prediction and nearly nullifies the ability for visual representation. While motion prediction improves, likely due to the unified task formulation, the overall performance degradation makes this latent space unsuitable. Forcing representation learning with this pixel-level latent space not only yields modest representation performance but also significantly weakens the other two tasks compared to learning in continuous semantic features. This suggests that high-level semantics are more effective than low-level details for understanding world states and distinguishing different contents in adjacent timesteps.
@@ 5 Conclusion and Future Work
In this paper, we presented EgoAgent, a joint predictive agent model that simultaneously learns to represent world observations, predict future world states, and generate 3D human motions in egocentric worlds. EgoAgent integrates these three abilities within a joint embedding-action-prediction architecture. Through extensive experiments, we demonstrated that EgoAgent outperforms existing state-of-the-art methods, with the joint learning of perception, prediction, and action tasks mutually reinforcing each other, leading to improved performance across all three capabilities. Future work will explore incorporating finer-grained hand representations for precise tasks such as object manipulation and integrating long-term memory mechanisms for tasks requiring extended temporal dependencies.

### v3

@@ 3.3 Task-Specific Inference
After training, EgoAgent inherently acquires the abilities to solve the three tasks designed to mirror human cognition:
World state prediction. Given a sequence of egocentric video frames and human motion, EgoAgent is capable of predicting the future world state $S_{t+1}$ by $S_{t+1}=\Theta(I_{1},A_{1},...,I_{t},A_{t}),$ where $\Theta$ denotes the EgoAgent model. Once predicted, the future world image can then be retrieved by measuring similarity within the learned feature space, enabling applications that require foresight into upcoming visual states.
Human action prediction. From an egocentric view and informed by previous body motions, EgoAgent can generate plausible future action $A_{t}$ represented as 3D human motion by $A_{t}=\Theta(I_{1},A_{1},...,A_{t-1},I_{t}).$ This predictive capability of human motion is essential for applications in humanoid robotics, virtual reality, and interactive gaming, where anticipating human-like movements is crucial.
Visual representation. EgoAgent learns world states as feature representations, allowing it to perceive meaningful semantic features from input images by $S_{t}=\Theta(I_{t}).$ These features can be directly applied to representation tasks such as image classification and visuomotor policy learning.
Table 1: Performance comparison across three tasks. (1) For world state prediction, we report the feature retrieval accuracy of the predicted next-frame features by the averaged Top1 accuracy and mAP on total timesteps $T=4$. (2) For human action prediction, we report the 3D motion prediction results on Ego-Exo4D by MPJPE (cm) and MPJVE (cm/s) with time gaps of 1/30 and 1/10 seconds. (3) For visual representation, we report the $k$-NN evaluation results on ImageNet-100 and ImageNet-1K by Top1 and Top5 accuracy (%). We evaluate with the officially released checkpoints of representation models trained on egocentric video datasets and train all motion prediction models from scratch. denotes the model trained by ourselves for a fair comparison, $\downarrow$ indicates lower is better.
@@ 4.1 Experimental Setups
Datasets. We train EgoAgent on two egocentric datasets, WalkingTours (WT) [44] and Ego-Exo4D [20]. WT is a video-only dataset with approximately 1.5M high-resolution frames. Ego-Exo4D has 221.26 hours of egocentric videos along with 376K manually labeled and 9.2M automatically generated 3D body poses. To eliminate the camera projection difference between Ego-Exo4D and WT, we undistort the raw fisheye sensor data in Ego-Exo4D into a pinhole camera model. We apply a sliding window filter of 20 frames to extract continuous pose sequences, resulting in 1,410,119 clips of synchronized egocentric videos and 3D body poses. Based on the EgoPose split in Ego-Exo4D, we reserve the validation set of Ego-Exo4D-v2333The whole Ego-Exo4D dataset is divided into Ego-Exo4D-v1 and Ego-Exo4D-v2 on the official website (https://ego-exo4d-data.org). for evaluation and use the remaining clips for training, yielding 1,378,672 training clips and 31,447 evaluation clips.
Implementation Details. To alleviate the training burden, we sample one image frame every five frames while retaining all 3D body poses. Specifically, for each clip with 20 frames, we divide it into $T=4$ time steps and sample the first image frame at each step. This results in every time step consisting of one image frame paired with five frames of 3D body poses. For self-supervised learning on egocentric videos, We follow [44] to avoid cropping noisy positive pairs and adopt two global crops, six local crops, and the same augmentations consistent with DINO [10].
Following the configurations in [24], we train EgoAgent with two model sizes, 300M and 1B, on the processed dataset for 72,000 iterations with a linear warm-up of 1800 iterations. We set the base learning rate to $6\times 10^{-4}$ with a cosine decay scheduler. The loss weights are set as $\lambda_{rep}=2,\lambda_{pred}=1,\lambda_{act}=3$, respectively. All models are trained from scratch using FP16 to speed up. The training takes 25/60 hours in total with a batch size of 1920 on 32/48 NVIDIA A100 GPUs for the 300M/1B model, respectively.
Figure 3: Retrieval results for egocentric world state prediction. Images correctly retrieved are highlighted with green borders.
@@ 4.2 World State Prediction
To evaluate the effectiveness of EgoAgent in predicting future states of the world, we perform an egocentric video prediction task with feature retrieval. Specifically, given the egocentric video frames and 3D human poses at a certain time step, we build a query set that stores the EgoAgent-predicted features of the next frame and a gallery set that contains the features directly extracted from all video frames. At time step $t$, if the predicted world state $S^{\prime}_{t+1}$ in the query set can correctly retrieve the corresponding $S_{t+1}$ in the gallery set, we treat it as a successful prediction. For traditional representation models, since there are no action conditions, the output state $S_{t+1}^{\prime}$ would be exactly the state at time step $t$, i.e., $S_{t}$. To avoid retrieving this exact same state, when querying with input image $I_{t}$, we ignore the scene state $S_{t}$ in the gallery set. Note that $S_{t}$ is NOT removed when testing EgoAgent. Following common practice in retrieval tasks, we use Top1 accuracy and mean average precision (mAP) as metrics.
As shown in Tab. 1, EgoAgent outperforms traditional representation models that trained with time-contrastive loss [32, 38, 44] or view-contrastive loss [10] on egocentric videos by a large margin. Specifically, EgoAgent-300M outperforms DoRA by +12.86% in Top1 accuracy and +13.05% in mAP, indicating that its predicted features are beyond the input images. In contrast, traditional models rely solely on semantic similarity, struggling to distinguish similar frames in a video. Furthermore, scaling up the model enhances performance, as EgoAgent-1B achieves further performance gains upon EgoAgent-300M with +3.42% Top1 accuracy and +3.90% mAP. Note that when trained on our combined dataset, DoRA exhibits performance degradation in both world state prediction and visual representation tasks. We attribute this to its tracker’s inability to track abrupt object motions in Ego-Exo4D. For further analysis, please refer to Sec. 4.4. Fig. 3 visualizes the retrieval results of predicted features, where DoRA fails to retrieve the correct future states when humans have large movements, while EgoAgent successfully predicts the future features that retrieve the correct next-frame images.
@@
To further evaluate whether EgoAgent accurately models the causal relationship between actions and future world states, we conduct a qualitative analysis by conditioning next-state predictions on various poses with the same initial observation, including dribbling, raising, shooting, and passing the ball. Fig. 4 shows the retrieved images, where all predicted features successfully retrieve images that accurately reflect the intended motion dynamics. This demonstrates EgoAgent’s ability to understand both the semantics of actions and their consequences in the egocentric world.
Figure 5: Visualizations of the 3D human motion prediction task. After observing the egocentric frames, EgoAgent can generate accurate human skeletons even though most of the body joints are not visible in the input frame.
@@
Quantitatively, as shown in Tab. 1, EgoAgent achieves state-of-the-art performance with the lowest MPJPE and competitive MPJVE. Specifically, EgoAgent-300M improves siMLPe [21] by -0.41 and -0.31 MPJPE on the evaluation set of 30 fps and 10 fps, respectively. When scaling up to 1B, the prediction errors are further reduced. These results demonstrate that EgoAgent can predict future actions accurately given egocentric observations and past actions.
We also evaluate the motion prediction results qualitatively. As illustrated in Fig. 5, EgoAgent produces relatively small errors compared to the ground truth. Compared to Diffusion Policy [13], EgoAgent achieves higher accuracy on non-visible joints. Specifically, in the early frames ($t_{1}$ and $t_{2}$), the missing right hand in the image leads to large prediction errors in Diffusion Policy, whereas EgoAgent accurately estimates hand positions by leveraging prior motion cues. Compared to HumanMAC [11], EgoAgent shows smaller accelerated errors. The predicted skeletons in HumanMAC gradually lean toward the floor ($t_{[3:5]}$) after several frames, whereas EgoAgent integrates visible body part information from egocentric observations to correct accelerated errors, leading to more stable motion predictions.
@@ 4.4 Visual Representation
We examine EgoAgent’s visual perception ability by two aspects: (1) image classification accuracy by $k$-NN on ImageNet-100 and ImageNet-1K [15], which is the widely-used golden standard for evaluating visual representations, and (2) egocentric manipulation success rate in an embodied simulator called TriFinger [48], which measures the transferability of learned visual representations to visual-motor control. For both tasks, the models are frozen and used to extract visual features. For image classification, we utilize a $k$-nearest neighbor classifier with $k=20$.
@@ 4.5 Ablation Study
Joint learning of representation, prediction, and action. EgoAgent demonstrates promising performance across the three tasks. To understand how these tasks interact during training, we conduct leave-one-out experiments and one-task-only experiments, as shown in Tab. 3 (a-g). We observe several key findings: (1) Removing any one task degrades the performance of the other two, highlighting the complementary roles of these tasks in learning. Specifically, removing the representation loss causes the largest drop in state prediction (-11.87% in Top1 accuracy and -13.55% in mAP), while removing the prediction loss significantly weakens representation learning, reducing Top1 accuracy by -2.52% on ImageNet-100 and -1.31% on ImageNet-1K. This suggests that state prediction and visual representation are closely interconnected. For motion prediction, excluding the prediction loss leads to the most significant decline, increasing MPJVE by 1.01. (2) Training on individual tasks yields worse performances on each task than learning all three tasks together, except for MPJPE in motion prediction. We suspect that when trained only on motion prediction, the model focuses excessively on minimizing joint position errors. However, without context from the other tasks, it struggles to maintain smooth, consistent motion across frames, resulting in a higher MPJVE of 2.52.
@@
Analysis of task dependencies. We also identify several interesting dependencies among the three tasks: (1) Representation serves as the foundation for both future state and action prediction as it captures the semantic information of the current world. (2) Prediction and action jointly provide effective supervision for representation learning by modeling world dynamics and predicting actions, improving representation by 1.16% in Top1 accuracy. However, training on prediction or action alone cannot enhance representation, as actions without accurate next-state predictions or next-state predictions conditioned on incorrect actions may provide noisy supervision. (3) Representation and action guide the model to integrate learned representations with action dynamics, leading to more accurate predictions of future states. Meanwhile, representation and prediction work together to help the model understand how actions influence state transitions, refining its ability to generate actions.
@@
Training in semantic feature space. To evaluate the effectiveness of training agent models with a high-level semantic feature space, we use a pretrained VQGAN [17], which is designed to reconstruct pixel-level details of images, as the image tokenizer. The results are shown in Tab. 3 (a,h,i). Training EgoAgent in VQGAN’s pretrained pixel-level latent space results in significant performance drops in world state prediction and nearly nullifies the ability for visual representation. While motion prediction improves, likely due to the unified task formulation, the overall performance degradation makes this latent space unsuitable. Forcing representation learning with this pixel-level latent space not only yields modest representation performance but also significantly weakens the other two tasks compared to learning in continuous semantic features. This suggests that high-level semantics are more effective than low-level details for understanding world states and distinguishing different contents in adjacent time steps.
@@ 5 Conclusion and Future Work
In this paper, we presented EgoAgent, a joint predictive agent model that simultaneously learns to represent world observations, predict future world states, and generate 3D human motions in egocentric worlds. EgoAgent integrates these three abilities within a joint embedding-action-prediction architecture with temporally asymmetric predictor-observer branches. Through extensive experiments, we demonstrated that EgoAgent outperforms existing state-of-the-art methods, with the joint learning of perception, prediction, and action tasks mutually reinforcing each other, leading to improved performance across all three capabilities. Future work will explore incorporating finer-grained hand representations for precise tasks such as object manipulation and integrating long-term memory mechanisms for tasks requiring extended temporal dependencies.
Acknowledgments: This work was partially supported by the Zhejiang Provincial Natural Science Foundation of China (No. LR25F020003), NSFC (No. 62402427, No. U24B20154), Zhejiang University Education Foundation Qizhen Scholar Foundation, and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University. We also acknowledge the support from the JC STEM Lab of AI for Science and Engineering, funded by the Hong Kong Jockey Club Charities Trust, and the Research Grants Council of Hong Kong (No. CUHK14213224).